# Variational Rank Reduction Autoencoders

### v1

@@ Abstract
Deterministic Rank Reduction Autoencoders (RRAEs) enforce by construction a regularization on the latent space by applying a truncated SVD. While this regularization makes Autoencoders more powerful, using them for generative purposes is counter-intuitive due to their deterministic nature. On the other hand, Variational Autoencoders (VAEs) are well known for their generative abilities by learning a probabilistic latent space. In this paper, we present Variational Rank Reduction Autoencoders (VRRAEs), a model that leverages the advantages of both RRAEs and VAEs. Our claims and results show that when carefully sampling the latent space of RRAEs and further regularizing with the Kullback-Leibler (KL) divergence (similarly to VAEs), VRRAEs outperform RRAEs and VAEs. Additionally, we show that the regularization induced by the SVD not only makes VRRAEs better generators than VAEs, but also reduces the possibility of posterior collapse. Our results include a synthetic dataset of a small size that showcases the robustness of VRRAEs against collapse, and three real-world datasets; the MNIST, CelebA, and CIFAR-10, over which VRRAEs are shown to outperform both VAEs and RRAEs on many random generation and interpolation tasks based on the FID score.
@@ 1 Introduction
Many real-life processes are probabilistic. Researchers, for decades, have been highlighting the variational aspect of the world planck1900zur ; wetterich2020probabilistic . Consequently, if we are to reproduce a certain process, probabilistic models are appealing. When learning to reproduce the probability distribution of a process, we don’t only learn a simple “input/output” relationship; we are also capable of replacing the process with the model, thus generating new samples.
More precisely, let $X$ be a set of training samples, following a certain distribution $a(X)$. Let us assume we have a model that can generate new samples with a distribution $b(X)$. We aim to encourage $a(X)$ and $b(X)$ to be as close as possible. Several techniques have been previously used for this type of generative modeling, but with various limitations. For instance, Gaussian Mixture Models (GMMs) reynolds2009gaussian ; mcnicholas2008parsimonious made strong assumptions about the structure of the data. Other techniques, such as Bayesian Networks pearl1988probabilistic ; govan2018bayesiannetwork ; govan2023bayesiannetwork or Hidden Markov Models rabiner1986introduction ; harte2021hiddenmarkov ; bourlard1994hidden , require an extensive inference procedure based on Markov Chain Monte Carlo (MCMC) gelman1996markov ; tierney1994markov .
On the other hand, as Neural Networks gained popularity due to their nonlinear capacities and speed during evaluation, Variational Autoencoders (VAEs) kingma2013autoencoding were introduced to learn probabilistic models while avoiding the limitations mentioned above. The goal of VAEs is to have a model that uses latent variables to sample the data. VAEs maximize the log likelihood of generating the training data by minimizing two objective functions,
@@
The first term maximizes the probability of the training data conditional on the latent variables, while the second term (called Kullback-Leibler (KL) divergence) enforces the predicted latent distribution to be as close as possible to a chosen distribution (usually standard normal), called the true prior doersch2021tutorialvariationalautoencoders ; DBLP:journals/corr/abs-1906-02691 .
The generative abilities of VAEs have made them very popular among probabilistic models. They have been used in many applications, including in the medical laptev2021generative , chemical lovric2021should , and engineering yan2020chiller ; regenwetter2022deep fields. VAEs are also the basic construction block for more complex models. For instance, some papers included discriminators gao2020zero ; niu2020lstm , others used regularization techniques wu2020vector ; hadjeres2017glsr , or replaced the KL divergence with an other regularizing term tolstikhin2019wassersteinautoencoders . While VAEs inherit the approximation abilities of Neural Networks and are fast during inference, they face two main limitations:
@@
    The KL divergence is the only regularization applied in a VAE. Accordingly, minimizing this term in some cases becomes crucial to obtain a meaningful latent space. Since during optimization we are trying to enforce both the reconstruction quality and the KL divergence, VAEs usually end up with worse reconstructions (i.e., blurrier images, in the case of image outputs) bredell2023explicitlyminimizingblurerror ; dalal2024shorttimefouriertransformdeblurring ; khan2018adversarialtrainingvariationalautoencoders .
@@
    Posterior collapse he2019lagginginferencenetworksposterior ; havrylov2020preventingposteriorcollapselevenshtein ; dang2024vanillavariationalautoencodersdetecting , which happens especially when the decoder is too large for the given amount of data. In this case, the probability distribution predicted by the encoder collapses to the true prior, thus deteriorating the generative abilities of VAEs.
On the other hand, recently, Rank Reduction Autoencoders (RRAEs) mounayer2025rankreductionautoencoders have been presented as a more stable version of Vanilla AEs. The main idea behind RRAEs is to include a truncated SVD in the latent space during training, enforcing the bottleneck by reducing the rank of the latent matrix instead of its dimension. Despite their deterministic nature, RRAEs have been shown to outperform other regularizing Autoencoders on tasks such as interpolation and random generation in their latent space. This is mainly due to the regularizing characteristics of the truncated SVD, without adding any terms to optimize in the loss. Yet, as RRAEs are not probabilistic models, they do not learn a distribution and therefore are less suitable to be used as generative models.
In this paper, we introduce Variational Rank Reduction Autoencoders (VRRAEs), a probabilistic version of RRAEs that learns a generative distribution of the data while benefiting from the regularization of the truncated SVD. We show throughout the paper, through conceptual and empirical evidence, that VRRAEs help mitigate both VAE’s limitations mentioned above (i.e., further regularization and robustness to posterior collapse). The resulting architecture produces clearer images and outperforms RRAEs and VAEs on MNIST, CelebA, and CIFAR-10 on almost all interpolation and random generation tasks. Furthermore, we revisit the dataset proposed by mounayer2025rankreductionautoencoders , but with a much smaller size (100 training images in total). We show that, on such a small dataset with local behavior, a typical VAE struggles to generate new samples due to posterior collapse, while VRRAEs overcome this limitation.
2 Proposed Method
We begin by defining the necessary notation and use those to present both VAEs and RRAEs.
@@ 2 Proposed Method
The objective of VAEs is to maximize the likelihood of the training data, which can be done by minimizing the following loss function kingma2013autoencoding ,
$\mathcal{L}_{VAE}=\frac{1}{N}\sum_{j=1}^{N}\left(\,\mathbb{E}_{q(\tilde{\alpha}\mid X_{j})}\left[-\log p(X_{j}\mid\tilde{\alpha})\right]\,+\,\mathrm{KL}\left(q(\tilde{\alpha}\mid X_{j})\,\|\,p(\tilde{\alpha})\right)\,\right),$ (1)
where $p(\tilde{\alpha})$ is a true prior, to which we want our latent distribution to resemble. When the prior is chosen to be standard normal with $p(\tilde{\alpha})=\mathcal{N}(0,I)$, and the decoder outputs a deterministic output $\tilde{X}$, the loss in Equation (1) can be re-written as kingma2013autoencoding ,
@@
where $\mathbf{1}_{k^{*}\times N}$ is a matrix of ones of dimension $k^{*}\times N$, $\odot$ is the Hadamard product (or element-wise multiplication), $\beta$ is a constant that scales the variance of the output distribution doersch2021tutorialvariationalautoencoders , $\|\cdot\|_{2}$ is the $L2$ norm, and “sum” indicates the sum of all of the entries of a matrix. Note that we choose to use the $L2$ norm for the reconstruction error throughout the paper, even though any other norm could have been used as well.
@@
On the other hand, Variational Autoencoders present two main limitations. First, since the KL divergence is the only regularization, and the loss to optimize combines both the reconstruction error and the KL divergence, the reconstruction error in some cases has to remain high if the space is to be regularized enough. Second, if the decoder is too large with respect to the complexity or size of the dataset, the training may steer the encoder to satisfy $q(\tilde{\alpha}|X_{j})\approx p(\tilde{\alpha})$, a phenomenon known as posterior collapse. While this reduces the magnitude of the KL loss term, it also causes the latent representation to lose its dependence on the input $X_{j}$, leading to an uninformative latent space and an increased reconstruction error.
@@
Note that during inference, the SVD is replaced by the projection over the basis found during training $\bar{U}_{f}$ (refer to mounayer2025rankreductionautoencoders for more details). In RRAEs, the bottleneck is not in the dimension of the latent space $\bar{Y}$, but in the rank of the truncated latent matrix, illustrated by the coefficients $\bar{\alpha}=\bar{S}\bar{V}^{T}$. Accordingly, both encoding/decoding maps $E$ and $D$ do not depend on the choice of the bottleneck $k^{*}$. Since the bottleneck is represented by $\bar{\alpha}$, it inherits some of the properties of the truncated singular values $\bar{S}$ and the truncated right singular vectors $\bar{V}^{T}$, mainly:
@@
The regularization in RRAEs has been shown to improve the behavior of autoencoders in mounayer2025rankreductionautoencoders . However, RRAEs are deterministic Autoencoders, which makes them less suitable for generating new samples in the latent space.
@@
Variational Rank Reduction Autoencoders (VRRAEs): In this part, we present the core of the paper, VRRAEs, an architecture designed to benefit from the advantages of both RRAEs and VAEs. Note that since the bottleneck in RRAEs is represented by $\bar{\alpha}$, only the truncated coefficients $\bar{\alpha}$ are sampled instead of sampling the reconstructed latent space $\bar{Y}$. As can be seen in the bottom of Figure 1, the SVD coefficients are sampled, similarly to how the bottleneck is sampled in VAEs (with a main difference discussed just after in the remark). The resulting, probabilistic, SVD coefficients $\tilde{\alpha}$ are then multiplied by the deterministic basis previously found $\bar{U}$ before being passed to the decoder.
Remark 1.
A main difference between VRRAEs and VAEs is that the means are found using $f=I$, the identity map. In other words, the expected values values $\mathbb{E}(\tilde{\alpha}_{i,j})=\bar{\alpha}_{i,j}$, $\forall i,j$, are the coefficients found by the SVD of $Y$. As detailed in Appendix A.1, and shown empirically in the ablation study in Section 4, this is crucial to preserve the properties and advantages of RRAEs.
@@
Further, note the VRRAEs, similarly to VAEs, regularize their latent space with the KL divergence. In practice, the same loss is implemented for both VAE and VRRAEs. Yet, note that when choosing the true prior to be $p(\tilde{\alpha})=\mathcal{N}(0,I)$, and by leveraging the properties of the SVD coefficients and the fact that $f=I$, the KL divergence for VRRAEs can be expressed as [^1] A detailed proof can be found in Appendix A.2.,
@@
where we used the same notation as Equation (2), and $\bar{S}$ is the truncated matrix containing the singular values on its diagonal, as defined in Equation (3). The term representing the mean, $(\text{diag}(\bar{S}))^{2}$, does not depend on the sample $j$, and simply consists of the square of the singular values. In other words, and as shown empirically in Appendix A.2, VRRAEs not only enforce a low rank of the latent matrix, but also encourage the singular values to be bounded. Though, just like VAEs, one has to find a compromise between bounding the singular values (minimizing the KL divergence), and reducing the reconstruction error (further details about the process of choosing $\beta$ for both VAEs, and VRRAEs can be found in Appendix A.3).
@@ 3 Results
This section presents empirical results to support the claims made when presenting VRRAEs. Throughout, we compare VRRAEs to three other models: a baseline diabolo AE, a Rank Reduction Autoencoder, and a Variational Autoencoder. We do not compare with other deterministic models since RRAEs outperformed them in mounayer2025rankreductionautoencoders . We also don’t compare with extensions of VAEs (e.g., discriminators or other regularizers than the KL-divergence), as these changes could be similarly applied to VRRAEs to improve their performance.
For training, all AE models share the same base architecture (i.e., encoder and decoder), which is detailed in Appendix A.3. The main differences between the models are the size of the latent space, set to $k^{*}$ for the Diabolo AE and the VAE, and to $L$ for the RRAE and the VRRAE (with a bottleneck of $k^{*}$ enforced via truncated SVD). Further, the optimal values of $\beta$ scaling the KL divergence are found by testing multiple values for both VAEs and VRRAEs and choosing the best model. The chosen values of $k^{*}$, $L$, and the process of choosing $\beta$ are given in Appendix A.3.
@@ 3.1 Synthetic Dataset
In this section, we test the Autoencoders discussed above on a challenging synthetic dataset to showcase the robustness of VRRAEs to posterior collapse. We chose the same dataset presented in mounayer2025rankreductionautoencoders , which consists of Gaussian bumps with a fixed spread and magnitude, on a 2D grid, that are moved on both the $x$ and $y$ axes. However, we reduced the number of training samples from $600$ to only $100$ (i.e., $10$ grid points in both $x$ and $y$). The small size of the dataset, as well as the challenging local behavior, makes it harder for autoencoders to find meaningful latent spaces without either collapsing the posterior or overfitting. To test the robustness of the AE architectures, we trained all four models 5 times with different seeds for initializing the network. We then randomly generated samples from the latent space of each Autoencoder by using a Gaussian Mixture Model. We quantified how well the generated images were by fitting a Gaussian to each new sample and finding the relative difference between the image and the fitted Gaussian with the right magnitude. The mean and standard deviation of the relative errors for random generation, as well as the reconstruction errors over a test set of $10000$ randomly chosen Gaussian curves can be found in Table 1.
The large standard deviation for VAEs shows that they might collapse their posterior depending on the seed. To further investigate wether it is posterior collapse, we visualize the expected value of the predicted posterior of each sample for a selected seed, in both VAEs and VRRAEs, by scattering them on a 2D plot in Figure 2. The small values of the means on the $y$-axis for VAEs shows that the second dimension of the latent space collapsed. On the other hand, we plot in red the latent space for many test gaussian images that represent a diagonal motion in the real space. As can be seen in the latent space of VRRAEs (Figure 2), the diagonal motion remains a diagonal motion in the latent space, which shows the interprability and meaningfulness of the latent space of VRRAEs.
Figure 2: Visualization of the second latent mean plotted against the first one for a certain seed in the latent space for VAEs (left) and VRRAEs (right). The collapse of the second latent dimension in VAEs is evidenced by the small variations along the $y$-axis (i.e., $0.02\leq\zeta_{i}\leq 0.09$ in Equation (4)). In contrast, VRRAEs do not suffer from collapse, and preserve the structure of a diagonal motion (highlighted in red), maintaining the interpretability of the latent representation.

### v2

@@ Abstract
Deterministic Rank Reduction Autoencoders (RRAEs) enforce by construction a regularization on the latent space by applying a truncated SVD. While this regularization makes Autoencoders more powerful, using them for generative purposes is counter-intuitive due to their deterministic nature. On the other hand, Variational Autoencoders (VAEs) are well known for their generative abilities by learning a probabilistic latent space. In this paper, we present Variational Rank Reduction Autoencoders (VRRAEs), a model that leverages the advantages of both RRAEs and VAEs. Our claims and results show that when carefully sampling the latent space of RRAEs and further regularizing with the Kullback-Leibler (KL) divergence (similarly to VAEs), VRRAEs outperform RRAEs and VAEs. Additionally, we show that the regularization induced by the SVD not only makes VRRAEs better generators than VAEs, but also reduces the possibility of posterior collapse. Our results include a synthetic dataset of a small size that showcases the robustness of VRRAEs against collapse, and three real-world datasets; the MNIST, CelebA, and CIFAR-10, over which VRRAEs are shown to outperform both VAEs and RRAEs on many random generation and interpolation tasks based on the FID score. We developed an open-source implementation of VRRAEs in JAX [1] (Equinox[2]), available at https://github.com/JadM133/RRAEs.git.
keywords:
Autoencoders , Latent space regularization , Generative models , Variational Autoencoders , Rank Reduction Autoencoders , Posterior collapse , Regularization.
journal: Neural Networks
[…]
inkscapelatex=false
[…]
[PIMM]organization=PIMM Laboratory, ENSAM,addressline=Blvd de l’Hôpital, city=Paris, postcode=75013, country=France
[…]
[Sara]organization=I3A, Universidad de Zaragoza,addressline=C. María de Luna, city=Zaragoza, postcode=50018, country=Spain
[…]
[SKF]organization=SKF Magnetic Mechatronics,addressline=Rue des Champs, city=Saint-Marcel, postcode=27950, country=France
[…]
[UNF]organization=University of North Florida,addressline=1 UNF Dr., city=Jacksonville, postcode=32224, country=United States
[…]
[CNRS]organization=CNRS@Create,addressline=1 CREATE Way, CREATE Tower, city=Singapore, postcode=138602, country=Singapore
@@ 1 Introduction
Many real-life processes are probabilistic. Researchers, for decades, have been highlighting the variational aspect of the world [3, 4]. Consequently, if we are to reproduce a certain process, probabilistic models are appealing. When learning to reproduce the probability distribution of a process, we don’t only learn a simple “input/output” relationship; we are also capable of replacing the process with the model, thus generating new samples.
More precisely, let $X$ be a set of training samples, following a certain distribution $a(X)$. Let us assume we have a model that can generate new samples with a distribution $b(X)$. We aim to encourage $a(X)$ and $b(X)$ to be as close as possible. Several techniques have been previously used for this type of generative modeling, but with various limitations. For instance, Gaussian Mixture Models (GMMs) [5, 6] made strong assumptions about the structure of the data. Other techniques, such as Bayesian Networks [7, 8, 9] or Hidden Markov Models [10, 11, 12], require an extensive inference procedure based on Markov Chain Monte Carlo (MCMC) [13, 14].
On the other hand, as Neural Networks gained popularity due to their nonlinear capacities and speed during evaluation, Variational Autoencoders (VAEs) [15] were introduced to learn probabilistic models while avoiding the limitations mentioned above. The goal of VAEs is to have a model that uses latent variables to sample the data. VAEs maximize the log likelihood of generating the training data by minimizing two objective functions,
@@
The first term maximizes the probability of the training data conditional on the latent variables, while the second term (called Kullback-Leibler (KL) divergence) enforces the predicted latent distribution to be as close as possible to a chosen distribution (usually standard normal), called the true prior [16, 17].
The generative abilities of VAEs have made them very popular among probabilistic models. They have been used in many applications, including in the medical [18], chemical [19], and engineering [20, 21] fields. VAEs are also the basic construction block for more complex models. For instance, some papers included discriminators [22, 23], others used regularization techniques [24, 25], or replaced the KL divergence with an other regularizing term [26]. While VAEs inherit the approximation abilities of Neural Networks and are fast during inference, they face two main limitations:
@@
    The KL divergence is the only regularization applied in a VAE. Accordingly, minimizing this term in some cases becomes crucial to obtain a meaningful latent space. Since during optimization we are trying to enforce both the reconstruction quality and the KL divergence, VAEs usually end up with worse reconstructions (i.e., blurrier images, in the case of image outputs) [27, 28, 29].
@@
    Posterior collapse [30, 31, 32], which happens especially when the decoder is too large for the given amount of data. In this case, the probability distribution predicted by the encoder collapses to the true prior, thus deteriorating the generative abilities of VAEs.
On the other hand, recently, Rank Reduction Autoencoders (RRAEs) [33] have been presented as a more stable version of Vanilla AEs. The main idea behind RRAEs is to include a truncated SVD in the latent space during training, enforcing the bottleneck by reducing the rank of the latent matrix instead of its dimension. Despite their deterministic nature, RRAEs have been shown to outperform other regularizing Autoencoders on tasks such as interpolation and random generation in their latent space. This is mainly due to the regularizing characteristics of the truncated SVD, without adding any terms to optimize in the loss. Yet, as RRAEs are not probabilistic models, they do not learn a distribution and therefore are less suitable to be used as generative models.
In this paper, we introduce Variational Rank Reduction Autoencoders (VRRAEs), a probabilistic version of RRAEs that learns a generative distribution of the data while benefiting from the regularization of the truncated SVD. We show throughout the paper, through conceptual and empirical evidence, that VRRAEs help mitigate both VAE’s limitations mentioned above (i.e., further regularization and robustness to posterior collapse). The resulting architecture produces clearer images and outperforms RRAEs and VAEs on MNIST, CelebA, and CIFAR-10 on almost all interpolation and random generation tasks. Furthermore, we revisit the dataset proposed by [33], but with a much smaller size (100 training images in total). We show that, on such a small dataset with local behavior, a typical VAE struggles to generate new samples due to posterior collapse, while VRRAEs overcome this limitation.
2 Background
We begin by defining the necessary notation and use those to present VAEs and RRAEs, both crucial for defining our model later.
@@ 2 Background
The objective of VAEs is to maximize the likelihood of the training data, which can be done by minimizing the following loss function [15],
[…]
@@
where $\mathbf{1}_{k^{*}\times N}$ is a matrix of ones of dimension $(k^{*}\times N)$, $\odot$ is the Hadamard product (or element-wise multiplication), $\beta$ is a constant that scales the variance of the output distribution [16], $\|\cdot\|_{2}$ is the $L2$ norm, and “sum” indicates the sum of all of the entries of a matrix. Note that we choose to use the $L2$ norm for the reconstruction error throughout the paper, even though any other norm could have been used as well.
@@
[…]
@@
Note that during inference, the SVD is replaced by the projection over the basis found during training $\bar{U}_{f}$ (refer to [33] for more details). In RRAEs, the bottleneck is not in the dimension of the latent space $\bar{Y}$, but in the rank of the truncated latent matrix, illustrated by the coefficients $\bar{\alpha}=\bar{S}\bar{V}^{T}$. Accordingly, both encoding/decoding maps $E$ and $D$ do not depend on the choice of the bottleneck $k^{*}$. Since the bottleneck is represented by $\bar{\alpha}$, it inherits some of the properties of the truncated singular values $\bar{S}$ and the truncated right singular vectors $\bar{V}^{T}$, mainly:
@@
The regularization in RRAEs has been shown to improve the behavior of autoencoders in [33]. However, RRAEs are deterministic Autoencoders, which makes them less suitable for generating new samples in the latent space.
@@
3 Proposed Model
In this part, we present the core of the paper, Variational Rank Reduction Autoencoders (VRRAEs), an architecture designed to benefit from the advantages of both RRAEs and VAEs. Note that since the bottleneck in RRAEs is represented by $\bar{\alpha}$, only the truncated coefficients $\bar{\alpha}$ are sampled instead of sampling the reconstructed latent space $\bar{Y}$. As can be seen in the bottom of Figure 1, the SVD coefficients are sampled, similarly to how the bottleneck is sampled in VAEs (with a main difference discussed just after in the remark). The resulting, probabilistic, SVD coefficients $\tilde{\alpha}$ are then multiplied by the deterministic basis previously found $\bar{U}$ before being passed to the decoder.
Remark 1
A main difference between VRRAEs and VAEs is that the means are found using $f=I$, the identity map. In other words, the expected values values $\mathbb{E}(\tilde{\alpha}_{i,j})=\bar{\alpha}_{i,j}$, $\forall i,j$, are the coefficients found by the SVD of $Y$. As detailed in A.1, and shown empirically in the ablation study in Section 5, this is crucial to preserve the properties and advantages of RRAEs.
@@ 3 Proposed Model
Further, note the VRRAEs, similarly to VAEs, regularize their latent space with the KL divergence. In practice, the same loss is implemented for both VAE and VRRAEs. Yet, note that when choosing the true prior to be $p(\tilde{\alpha})=\mathcal{N}(0,I)$, and by leveraging the properties of the SVD coefficients and the fact that $f=I$, the KL divergence for VRRAEs can be expressed as [^1] A detailed proof can be found in A.2.,
@@
where we used the same notation as Equation (2), and $\bar{S}$ is the truncated matrix containing the singular values on its diagonal, as defined in Equation (3). The term representing the mean, $(\text{diag}(\bar{S}))^{2}$, does not depend on the sample $j$, and simply consists of the square of the singular values. In other words, and as shown empirically in A.2, VRRAEs not only enforce a low rank of the latent matrix, but also encourage the singular values to be bounded. Though, just like VAEs, one has to find a compromise between bounding the singular values (minimizing the KL divergence), and reducing the reconstruction error (further details about the process of choosing $\beta$ for both VAEs, and VRRAEs can be found in A.3).
@@ 4 Results
This section presents empirical results to support the claims made when presenting VRRAEs. Throughout, we compare VRRAEs to three other models: a baseline diabolo AE, a Rank Reduction Autoencoder, and a Variational Autoencoder. VRRAEs are also compared to other deterministic Autoencoders in A.6. We do not compare our model with extensions of VAEs (e.g., discriminators, annealing of $\beta$ values, or using other regularizers than the KL-divergence), as these changes could be similarly applied to VRRAEs to improve their performance.
For training, all AE models share the same base architecture (i.e., encoder and decoder), which is detailed in A.3. The main differences between the models are the size of the latent space, set to $k^{*}$ for the Diabolo AE and the VAE, and to $L$ for the RRAE and the VRRAE (with a bottleneck of $k^{*}$ enforced via truncated SVD). Further, the optimal values of $\beta$ scaling the KL divergence are found by testing multiple values for both VAEs and VRRAEs and choosing the best model. The chosen values of $k^{*}$, $L$, and the process of choosing $\beta$ are given in A.3.
Readers interested in more details about the training times, or the effect of the batch size can refer to both A.4, and A.5. Further, the code used to produce the results in this paper is part of an open-source library in JAX [1] (Equinox [2]), available at https://github.com/JadM133/RRAEs.git.
@@
Figure 2: Visualization of the second latent mean plotted against the first one for a certain seed in the latent space for VAEs (left) and VRRAEs (right). The collapse of the second latent dimension in VAEs is evidenced by the small variations along the $y$-axis (i.e., $0.02\leq\zeta_{i}\leq 0.09$ in Equation (4)). In contrast, VRRAEs do not suffer from collapse, and preserve the structure of a diagonal motion (highlighted in red), maintaining interpretability.
@@ 4.1 Synthetic Dataset
In this section, we test the Autoencoders discussed above on a challenging synthetic dataset to showcase the robustness of VRRAEs to posterior collapse. We chose the same dataset presented in [33], which consists of Gaussian bumps with a fixed spread and magnitude, on a 2D grid, that are moved along both the $x$ and $y$ axes. However, we reduced the number of training samples from $600$ to only $100$ (i.e., $10$ grid points in both $x$ and $y$). The small size of the dataset, as well as the challenging local behavior, makes it harder for autoencoders to find meaningful latent spaces without either collapsing the posterior or overfitting. To test the robustness of the AE architectures, we trained all four models 5 times with different seeds for initializing the network. We then randomly generated samples from the latent space of each Autoencoder by using a Gaussian Mixture Model. We quantified how well the generated images were by fitting a Gaussian to each new sample and finding the relative difference between the image and the fitted Gaussian with the right magnitude. The mean and standard deviation of the relative errors for random generation, as well as the reconstruction errors over a test set of $10000$ randomly chosen Gaussian curves can be found in Table 1.
The large standard deviation for VAEs shows that they might collapse their posterior depending on the seed. To further investigate wether it is posterior collapse, we visualize the expected value of the predicted posterior of each sample for a selected seed, in both VAEs and VRRAEs, by scattering them on a 2D plot in Figure 2. The small values of the means on the $y$-axis for VAEs shows that the second dimension of the latent space collapsed. On the other hand, we plot in red the latent space for many test gaussian images that represent a diagonal motion in the real space. As can be seen in the latent space of VRRAEs (Figure 2), the diagonal motion remains a diagonal motion in the latent space, which shows the interpretability of the latent space of VRRAEs.